Weights of Pink salmon are N(4.8, 0.4) pounds.
30% of Pink salmon weigh less than

Fishermen consider a fish to be a "lunker" if it is in the top 4% of size. How much must a Pink salmon weigh to be considered a lunker?
Any salmon over
pounds is considered a linker. (round to the nearest 10th)

Answer by Boreal(15235)

96%ile is 1.75 for z
z=(x-mean)/sd
1.75*0.4=x-4.8
0.7=x-4.8
x=5.5 pounds
The other parts aren't clear.